# From Lipschitz Embedding to Lipschitz Equivalence between Dust-like Self-similar Sets

### v1

@@ 5. Lipschitz equivalence via the homogeneity of $F$
with $\operatorname{Lip}(g_{k})\leq\operatorname{Lip}(\widetilde{f})\leq L$. Similar to the second paragraph of the proof of Lemma 4.7, we have $\psi_{\bm{j}_{k}\theta_{k}}^{-1}(A_{k,1})\in\mathcal{K}$ for all $k$. Then, combining Lemma 2.2, Remark 2.3 and Lemma 5.4(2), we obtain a Lipschitz surjection $g:E\to F$ for some $E\in\mathcal{K}$ such that $\operatorname{Lip}(g)\leq L$ and
@@ 6. A counterexample in the inhomogeneous case
To state a mass decomposition lemma in [22], write $\mathbb{Z}[r^{s}]$ be the ring generated by $r^{s}$ and the integer set $\mathbb{Z}$. With a slight abuse of notation, here we pick $\delta=r$ and let
$\left\{\begin{array}[]{l}I_{n}:=\{\bm{i}\in I^{*}:r_{\bm{i}}\leq\delta^{n}<r_{\bm{i}^{-}}\},\\ J_{n}:=\{\bm{j}\in J^{*}:\lambda_{\bm{j}}\leq\delta^{n}<\lambda_{\bm{j}^{-}}\}.\end{array}\right$.
@@
In this example, the ratio root $r$ equals $3^{-1}$. So $\delta=3^{-1}$ and $J_{1}=J$. Write
@@
For such parameters $r$ and $\Omega$, there exists a large integer $c$ as specified in Lemma 6.1. We claim that for every $k\geq 1$, there exists a partition $I_{1+kc}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\tilde{I}_{\bm{j}}$ such that $\sum_{\bm{i}\in\tilde{I}_{\bm{j}}}r_{\bm{i}}^{s}=\lambda_{\bm{j}}^{s}$ for each $\bm{j}\in J_{1+(k-1)c}$.
The claim can be proved by induction. Clearly, picking $c$ large enough at the beginning, every cell in $\{K_{\bm{i}}:\bm{i}\in I^{3}\}$ is a finite union of cells in $\{K_{\omega}:\omega\in I_{1+c}\}$. So we can transform the previous partition of $I^{3}$ into another partition $I_{1+c}=\bigcup_{\bm{j}\in J_{1}}\tilde{I}_{\bm{j}}$ with $\sum_{\bm{i}\in\tilde{I}_{\bm{j}}}r_{\bm{i}}^{s}=\lambda_{\bm{j}}^{s}$ for each $\bm{j}\in J_{1}$. This settles the case $k=1$.
@@
Let $\bm{j}\in J_{1+(k-1)c}$. Note that by the induction hypothesis, $\bm{j}$ corresponds to some $\tilde{I}_{\bm{j}}\subset I_{1+kc}$ with
@@
where $\bm{j}\prec\eta$ means that $\bm{j}$ is a prefix of $\eta$. In other words, writing $\bigcup_{\bm{i}\in\tilde{I}_{\bm{j}}}K_{\bm{i}}, EE$ is simply an element in $\mathcal{B}_{1+kc}$ (as in Lemma 5.4) with $\mu(E)=\sum_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\lambda_{\eta}^{s}$, where each $\lambda_{\eta}^{s}$ is a member of $r^{(1+kc)s}\Omega$ (recall (6.1)). It then follows from Lemma 6.1 that there exists a partition $E=\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}E_{\eta}$, where every $E_{\eta}$ is a member of $\mathcal{B}_{1+(k+1)c}$ with $\mu(E_{\eta})=\lambda_{\eta}^{s}$. By definition, there is some $\tilde{I}_{\eta}\subset I_{1+(k+1)c}$ such that $E_{\eta}=\bigcup_{\kappa\in\tilde{I}_{\eta}}K_{\kappa}$. We conclude that
(6.2) $\bigcup_{\omega\in\tilde{I}_{\bm{j}}}K_{\omega}=E=\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}E_{\eta}=\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\bigcup_{\kappa\in\tilde{I}_{\eta}}K_{\kappa},$
@@
$\bigcup_{\kappa\in I_{1+(k+1)c}}K_{\kappa}=K$ $=\bigcup_{\omega\in I_{1+kc}}K_{\omega}$
$=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\bigcup_{\omega\in\tilde{I}_{\bm{j}}}K_{\omega}$ (since $I_{1+kc}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\tilde{I}_{\bm{j}}$)
$=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\bigcup_{\kappa\in\tilde{I}_{\eta}}K_{\kappa}$ (by (6.2))
$=\bigcup_{\eta\in J_{1+kc}}\bigcup_{\kappa\in\tilde{I}_{\eta}}K_{\kappa},$
@@
such that $\sum_{\kappa\in\tilde{I}_{\eta}}r_{\kappa}^{s}=\mu(E_{\eta})=\lambda_{\eta}^{s}$ for all $\eta\in J_{1+kc}$. This completes the induction.
The mass distribution in our claim naturally induces a Lipschitz surjection from $K$ to $F$ as follows. For $x\in K$ and $k\geq 1$, let $\bm{i}_{k}$ be the unique word in $I_{1+kc}$ such that $x\in K_{\bm{i}_{k}}$. Note that $\bm{i}_{1}\prec\bm{i}_{2}\prec\cdots$ and $\{x\}=\bigcap_{k}K_{\bm{i}_{k}}$. By our claim, there exists $\bm{j}_{k}\in J_{1+(k-1)c}$ such that $\bm{i}_{k}\in\tilde{I}_{\bm{j}_{k}}$. Moreover, it follows from (6.2) that
$K_{\bm{i}_{k+1}}\subset K_{\bm{i}_{k}}\subset\bigcup_{\omega\in\tilde{I}_{\bm{j}_{k}}}K_{\omega}=\bigcup_{\eta\in J_{1+kc}:\bm{j}_{k}\prec\eta}\bigcup_{\kappa\in\tilde{I}_{\eta}}K_{\kappa}.$
In other words, $\bm{i}_{k+1}\in\bigcup_{\eta\in J_{1+kc}:\bm{j}_{k}\prec\eta}\tilde{I}_{\eta}$. So $\bm{j}_{k}\prec\bm{j}_{k+1}$, that is, $\bm{j}_{1}\prec\bm{j}_{2}\prec\cdots$ as well. Then we just set $g(x)$ to be the single point in $\bigcap_{k}F_{\bm{j}_{k}}$. In this way, it is not hard to see that $g(K)=F$. It remains to prove that $g$ is Lipschitz.
Let $x,x^{\prime}\in K$ be distinct and let $k_{*}$ be the maximal integer such that $x,x^{\prime}\in K_{\omega_{*}}$ for a common $\omega_{*}\in I_{1+k_{*}c}$. So $|x-x^{\prime}|\approx\operatorname{diam}(K_{\omega_{*}})=r_{\omega_{*}}\approx 3^{-(1+k_{*}c)}$. Since $g$ sends $x,x^{\prime}$ into a specific cell in $\{F_{\bm{j}}:\bm{j}\in J_{1+(k_{*}-1)c}\}$, we have
@@
Acknowledgements. The research of the authors is partially supported by NSFC grant 12371089, and the Fundamental Research Funds for the Central Universities of China grant 2024FZZX02-01-01.

### v2

@@ 5. Lipschitz equivalence via the homogeneity of $F$
with $\operatorname{Lip}(g_{k})\leq\operatorname{Lip}(\widetilde{f})\leq L$. Similar to the fourth paragraph of the proof of Lemma 4.7, we have $\psi_{\bm{j}_{k}\theta_{k}}^{-1}(A_{k,1})\in\mathcal{K}$ for all $k$. Then, combining Lemma 2.2, Remark 2.3 and Lemma 5.4(1), we obtain a Lipschitz surjection $g:E\to F$ for some $E\in\mathcal{K}$ such that $\operatorname{Lip}(g)\leq L$ and
@@ 6. A counterexample in the inhomogeneous case
To state a mass decomposition lemma in [22], let $\mathbb{Z}[r^{s}]$ be the ring generated by $r^{s}$ and the integer set $\mathbb{Z}$. With a slight abuse of notation, here we pick $\delta=r$ and let
(6.1) $\left\{\begin{array}[]{l}I_{n}:=\{\bm{i}\in I^{*}:r_{\bm{i}}\leq\delta^{n}<r_{\bm{i}^{-}}\},\\ J_{n}:=\{\bm{j}\in J^{*}:\lambda_{\bm{j}}\leq\delta^{n}<\lambda_{\bm{j}^{-}}\}.\end{array}\right$.
@@
In this example, the ratio root $r$ equals $3^{-1}$, so $\delta=3^{-1}$ and $J_{1}=J$. Write
@@
For such parameters $r$ and $\Omega$, there exists a large integer $c$ as specified in Lemma 6.1. We claim that for every $k\geq 1$, there exists a partition $I_{1+kc}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\tilde{I}_{\bm{j}}$ such that $\sum_{\bm{i}\in\tilde{I}_{\bm{j}}}r_{\bm{i}}^{s}=\lambda_{\bm{j}}^{s}$ for each $\bm{j}\in J_{1+(k-1)c}$, where $I_{n}$ and $J_{n}$ are as defined in (6.1).
The claim is proved by induction. Clearly, picking $c$ large enough initially, every cell in $\{K_{\bm{i}}:\bm{i}\in I^{3}\}$ can be expressed as a finite union of cells in $\{K_{\omega}:\omega\in I_{1+c}\}$. Since $J_{1}=J$, we can transform the previous partition of $I^{3}$ into another partition $I_{1+c}=\bigcup_{\bm{j}\in J_{1}}\tilde{I}_{\bm{j}}$ with $\sum_{\bm{i}\in\tilde{I}_{\bm{j}}}r_{\bm{i}}^{s}=\lambda_{\bm{j}}^{s}$ for each $\bm{j}\in J_{1}$. This settles the case $k=1$.
@@
Let $\bm{j}\in J_{1+(k-1)c}$. By the induction hypothesis, $\bm{j}$ corresponds to some $\tilde{I}_{\bm{j}}\subset I_{1+kc}$ with
@@
where $\bm{j}\prec\eta$ means that $\bm{j}$ is a prefix of $\eta$. In other words, writing $E_{\bm{j}}:=\bigcup_{\bm{i}\in\tilde{I}_{\bm{j}}}K_{\bm{i}}$, $E_{\bm{j}}$ is simply an element in $\mathcal{B}_{1+kc}$ (as in Lemma 5.4) with $\mu(E_{\bm{j}})=\sum_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\lambda_{\eta}^{s}$, where each $\lambda_{\eta}^{s}$ is a member of $r^{(1+kc)s}\Omega$ (recall (6.2)). It then follows from Lemma 6.1 that there exists a partition $E_{\bm{j}}=\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}E_{\eta}$, where every $E_{\eta}$ is a member of $\mathcal{B}_{1+(k+1)c}$ with $\mu(E_{\eta})=\lambda_{\eta}^{s}$. By definition, there is some $\tilde{I}_{\eta}\subset I_{1+(k+1)c}$ such that $E_{\eta}=\bigcup_{\xi\in\tilde{I}_{\eta}}K_{\xi}$. We conclude that
(6.3) $\bigcup_{\bm{i}\in\tilde{I}_{\bm{j}}}K_{\bm{i}}=E_{\bm{j}}=\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}E_{\eta}=\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\bigcup_{\xi\in\tilde{I}_{\eta}}K_{\xi},$
@@
(6.4) $\bigcup_{\bm{j}\in J_{1+(k-1)c}}\bigcup_{\bm{i}\in\tilde{I}_{\bm{j}}}K_{\bm{i}}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\bigcup_{\eta\in J_{1+kc}:\bm{j}\prec\eta}\bigcup_{\xi\in\tilde{I}_{\eta}}K_{\xi}=\bigcup_{\eta\in J_{1+kc}}\bigcup_{\xi\in\tilde{I}_{\eta}}K_{\xi},$
because $J_{1+kc}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\{\eta\in J_{1+kc}:\bm{j}\prec\eta\}$. On the other hand, since $I_{1+kc}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\tilde{I}_{\bm{j}}$ (the induction hypothesis),
$\bigcup_{\bm{i}\in I_{1+kc}}K_{\bm{i}}=\bigcup_{\bm{j}\in J_{1+(k-1)c}}\bigcup_{\bm{i}\in\tilde{I}_{\bm{j}}}K_{\bm{i}}.$
Combining this with (6.4), we have
$\bigcup_{\xi\in I_{1+(k+1)c}}K_{\xi}=K=\bigcup_{\bm{i}\in I_{1+kc}}K_{\bm{i}}=\bigcup_{\eta\in J_{1+kc}}\bigcup_{\xi\in\tilde{I}_{\eta}}K_{\xi},$
@@
such that $\sum_{\xi\in\tilde{I}_{\eta}}r_{\xi}^{s}=\mu(E_{\eta})=\lambda_{\eta}^{s}$ for all $\eta\in J_{1+kc}$. This completes the induction.
The claim naturally induces a Lipschitz surjection from $K$ to $F$ as follows. For $x\in K$ and $k\geq 1$, let $\bm{i}_{k}$ be the unique word in $I_{1+kc}$ such that $x\in K_{\bm{i}_{k}}$. Note that $\bm{i}_{1}\prec\bm{i}_{2}\prec\cdots$ and $\{x\}=\bigcap_{k}K_{\bm{i}_{k}}$. By our claim, there exists $\bm{j}_{k}\in J_{1+(k-1)c}$ such that $\bm{i}_{k}\in\tilde{I}_{\bm{j}_{k}}$. Moreover,
(6.5) $K_{\bm{i}_{k+1}}\subset K_{\bm{i}_{k}}\subset\bigcup_{\omega\in\tilde{I}_{\bm{j}_{k}}}K_{\omega}=\bigcup_{\eta\in J_{1+kc}:\bm{j}_{k}\prec\eta}\bigcup_{\xi\in\tilde{I}_{\eta}}K_{\xi},$
where the last equality follows from (6.3). Since $\bm{i}_{k+1}\in I_{1+(k+1)c}$ and $\bigcup_{\eta\in J_{1+kc}:\bm{j}_{k}\prec\eta}\tilde{I}_{\eta}\subset I_{1+(k+1)c}$, (6.5) implies that
$\bm{i}_{k+1}\in\bigcup_{\eta\in J_{1+kc}:\bm{j}_{k}\prec\eta}\tilde{I}_{\eta}.$
Combining this with our choice of $\bm{j}_{k+1}$ (i.e., the word in $J_{1+kc}$ such that $\bm{i}_{k+1}\in\widetilde{I}_{{\bm{j}}_{k+1}}$), we see that $\bm{j}_{k+1}\in\{\eta\in J_{1+kc}:\bm{j}_{k}\prec\eta\}$. In particular, $\bm{j}_{k}\prec\bm{j}_{k+1}$. So $\bm{j}_{1}\prec\bm{j}_{2}\prec\cdots$ as well and we set $g(x)$ to be the single point in $\bigcap_{k}F_{\bm{j}_{k}}$.
To see the surjectivity of $g$, let $y\in F$ be arbitrary. For each $k\geq 1$, let $\xi_{k}$ be the unique word in $\widetilde{J}_{1+(k-1)c}$ such that $y\in F_{\xi_{k}}$. From the previous induction argument, it follows that $B_{k}:=\bigcup_{\bm{i}\in\widetilde{I}_{\xi_{k}}}K_{\bm{i}}$ is non-empty, compact and nested decreasing in $k$ (see (6.3)). Moreover, $g$ maps the non-empty intersection $\bigcap_{k=1}^{\infty}B_{k}$ to the singleton $\{y\}$. So $g(K)=F$.
It remains to prove that $g$ is Lipschitz. Let $x,x^{\prime}\in K$ be distinct and let $k_{*}$ be the maximal integer such that $x,x^{\prime}\in K_{\omega_{*}}$ for a common $\omega_{*}\in I_{1+k_{*}c}$. So $|x-x^{\prime}|\approx\operatorname{diam}(K_{\omega_{*}})=r_{\omega_{*}}\approx 3^{-(1+k_{*}c)}$. Since $g$ sends $x,x^{\prime}$ into a specific cell in $\{F_{\bm{j}}:\bm{j}\in J_{1+(k_{*}-1)c}\}$, we have
@@
Acknowledgements. The research of the authors is partially supported by NSFC grants 12371089, 12501113, 12531004, and the Fundamental Research Funds for the Central Universities of China grant 2024FZZX02-01-01.